# Orbit lengths for promotion on 2-row and near-hook tableaux

### v1

@@ 6.2 Quadratic orbit lengths for near-hooks
In the tableau $T[15]$ below, the singletons are 3, 9, and 15 (highlighted in orange), while the runs are $(14,4,5)$ and $(8,11)$, with the run $(14,4,5)$ highlighted in green and the run $(8,11)$ in yellow:
@@ 11 $13$ 33 44 55 88 99 $11$ $14$ $15$
    If the 3 (now a 2) was the first entry of the run, the $n$ is now a member of the run, and the $\ell-2$ is a singleton nested inside the run. Thus, the the run length stays the same, its gaps to other runs stay the same (since only something not in the first column changed), and the number of singletons stays the same. The singleton just decreased by 1 from $\ell-1$ to $\ell-2$, so the gaps on the singleton track also stay the same.
@@
Combining Lemmas 6.2.4 and 6.2.5 completes the proof of Theorem 6.2.2. ∎
@@ Acknowledgments
I’m thankful to Oliver Pechenik for getting me started on this project and for several helpful discussions about it. I’m also thankful to Jessica Striker for a helpful conversation related to this. I used Sage [Ste+25] to help formulate and test various conjectures.

### v2

@@ 6.2 Quadratic orbit lengths for near-hooks
In the tableau $T[16]$ below, the gaps are large enough that the runs are just the sequences of consecutive numbers not in the first row:
11 $16$ 44 $15$ 55 99 $10$ $11$
The runs are $(4,5)$ and $(9,10,11)$, highlighted in green and yellow, while 15 is a singleton, highlighted in orange. Thus, $r=2$ and $s=1$.
Example 6.2.2.
In the tableau $T[15]$ below, the runs and singletons are not quite what we might expect, since some singletons are nested inside runs. The singletons are 3, 9, and 15 (highlighted in orange), while the runs are $(14,4,5)$ and $(8,11)$, with the run $(14,4,5)$ highlighted in green and the run $(8,11)$ in yellow:
@@ 11 $13$ 33 44 55 88 99 $11$ $14$ $15$
    If the 3 (now a 2) was the first entry of the run, the $n$ is now a member of the run, and the $\ell-2$ is a singleton nested inside the run. Thus, the run length stays the same, its gaps to other runs stay the same (since only something not in the first column changed), and the number of singletons stays the same. The singleton just decreased by 1 from $\ell-1$ to $\ell-2$, so the gaps on the singleton track also stay the same.
@@
Combining Lemmas 6.2.5 and 6.2.6 completes the proof of Theorem 6.2.3. ∎
@@ Acknowledgments
I’m thankful to Oliver Pechenik for getting me started on this project and for several helpful discussions about it. I’m also thankful to Jessica Striker for a helpful conversation related to this. I used Sage [10] to help formulate and test various conjectures.